OpenStudy (kainui):

A fun integral question:

OpenStudy (kainui):

\[\LARGE \int\limits_0^\infty \frac{\tan^{-1}(ax)-\tan^{-1}(x)}{x}dx\]

hartnn (hartnn):

I am thinking Differentiation Under Integral Sign Technique

hartnn (hartnn):

\(\Large \dfrac{\partial I}{\partial a} = \int \limits_0^\infty \dfrac{1}{a^2 x^2+1 }dx\)

hartnn (hartnn):

I = pi/2 ln a

OpenStudy (kainui):

Interesting! I had a different path to the solution in mind. \[\Large \int\limits_0^\infty \int\limits _x^{ax}\frac{1}{x(1+y^2)}dydx\] Reverse the order of integration:\[\Large \int\limits_0^\infty \int\limits _{y/a}^y \frac{1}{x(1+y^2)}dxdy = \int\limits_0^\infty \frac{lny-lny+lna}{1+y^2}dy \] \[\Large \ln a \frac{\pi}{2}\] Same answer, but kind of like opposite methods haha.

hartnn (hartnn):

I am checking if Laplace Transform helps to find us an alternate way :)

hartnn (hartnn):

Interesting kainui, can you brief on how you got that first double integral ?

hartnn (hartnn):

y = arctan x ?

OpenStudy (kainui):

Yeah, try evaluating the integral for dy to get the original integral back.

hartnn (hartnn):

that path is easy, but reverse part is quite difficult

OpenStudy (kainui):

Since every integral ends with a function evaluated at two points and them subtracted, I just sort of decided to try that.

hartnn (hartnn):

@SithsAndGiggles and @ganeshie8 might also find this interesting :)

OpenStudy (kainui):

Similarly you can make a messed up definition of the derivative if you do something fun like this haha, \[\lim_{h \to 0} \int\limits_{f(x)} ^{f(x+h)} \frac{dx}{h}= f'(x)\] But at anyway, I can try to give more explanation or reasoning if I didn't really explain what you were asking, I'm not sure I answered your question.

hartnn (hartnn):

you did, clearly :)

OpenStudy (kainui):

@hartnn If you know any good problems to solve with differentiation under the integral sign I'd like to see one since I don't get enough practice with that since it's usually pretty hard to see when to use it.

hartnn (hartnn):

I generally think of DUIS when there is one constant in a complicated looking function to be integrated.

hartnn (hartnn):

I did one DUIS problem with ganeshie few days ago here, but my profile isn't loading that many questions of the past... . I'll try to find one from other site..

hartnn (hartnn):

wiki examples are quite good... http://en.wikipedia.org/wiki/Differentiation_under_the_integral_sign

OpenStudy (anonymous):

I've yet to get a grasp on the intuition/strategic mindset is DUIS... Any good resources or tutorials for that?